Find two consecutive integers such that the sum of 4 times the first integer and 8 times second integer is 68.
Answer by elima(1433)

Let x be the =first integer
x+1 = the second intger
4x+8(x+1)=68
4x+8x+8=68
12x+8=68
12X=60
X=5
The first integer = 5 and the second integer;
x+1=5+1=6
Hope you understand
=)